A sample of final exam scores is normally distributed with a mean equal to 29 and a variance equal to 16.

Part (a)

What percentage of scores are between 25 and 33? (Round your answer to two decimal places.)

Part (b)

What raw score is the cutoff for the top 10% of scores? (Round your answer to one decimal place.)

Part (c)

What is the proportion below 23? (Round your answer to four decimal places.)


Part (d)

What is the probability of a score less than 35? (Round your answer to four decimal places.)
Answer #1

Solution :

Given that ,

mean = \mu = 29

standard deviation = \sigma = \sqrt 16 = 4

a)P( 25< x < 33) = P[(25 - 29)/4 ) < (x - \mu ) /\sigma  < (33-29) /4 ) ]

= P( -1< z <1 )

= P(z <1 ) - P(z < -1)

Using standard normal table

= 0.8413 - 0.1587 = 0.6827

Probability = 0.6827 = 68.27%

b) top 10 % =0.10

P(Z > z ) = 0.10

1- P(z < z) =0.10

P(z < z) = 1-0.10 = 0.90

z = 1.28

Using z-score formula,

x = z * \sigma + \mu

x = 1.28* 4 + 29

x = 34.1

c)

P(x < 23 ) = P[(x - \mu ) / \sigma < (23 - 29) /4 ]

= P(z <-1.5 )

= 0.0668

proportion = 0.0668

d)

P(x < 35) = P[(x - \mu ) / \sigma < (35 - 29) /4 ]

= P(z <1.5 )

= 0.9332

probability = 0.9332